# Accelerating High-Order Finite Element Simulations at Extreme Scale with FP64 Tensor Cores

### 10 Mar 2026 v1

@@ I Introduction
Starting in 2020 with the Ampere A100 GPU, NVIDIA introduced FP64 tensor cores supporting double precision MMA (DMMA) instructions. For double precision calculations, these tensor cores deliver a 2 increase in peak floating-point rates [undefd]. In recent years, other vendors have introduced tensor cores, or analogous concepts, in their compute architectures. In this paper, we focus specifically on the programming of tensor cores on two of NVIDIA’s state-of-the-art chips: the Grace Hopper GH200 Superchip [undefe] and the Grace Blackwell GB200 Superchip [undefb].
@@
We build on this prior work and apply tensor cores to a full finite element production application—a digital twin for tsunami early warning [undefl], where the coupled acoustic–gravity wave propagation is solved as a first-order system of partial differential equations (PDEs) discretized with high-order finite elements. Specifically, we decompose the tensor contractions into small GEMMs of order and program the FP64 tensor cores directly in CUDA kernels. We show how directly programming tensor cores on state-of-the-art chips like NVIDIA’s Grace Hopper GH200 and Grace Blackwell GB200 can lead to significant performance and energy efficiency gains in a complex application code.
@@
To the best of our knowledge, the first example of using directly programmed FP64 tensor cores in a complex, PDE-based HPC application code, in order to improve throughput of high-order finite element kernels (by up to 59%).
@@
Demonstration of full-system scalability of the application code for each of the implemented kernel versions on a leading-edge supercomputer.
The rest of the paper is organized as follows. Section II provides a brief overview of the MFEM library, the finite element algorithms and application used, and the NVIDIA tensor cores. We then discuss how we optimized the key finite element kernels in Section III. Section IV shows the performance and energy efficiency improvements from our optimizations. In Section V and Section VI, respectively, we describe related work and summarize our conclusions.
@@ II-A MFEM library
II-B Finite elements assembly algorithms on GPUs
@@
GPU-accelerated PA kernels cover a broad range of operators, enabling end-to-end acceleration of matrix-free solvers through low-order-refined preconditioning [undeft]. Kernel fusion methods enhance strong scalability and allow attaining peak performance on problem sizes 5–10 times smaller than conventional methods. These matrix-free approaches, distinct from traditional sparse matrix assembly, significantly reduce memory footprint by avoiding global matrix storage, enhance arithmetic intensity, achieving near optimal floating-point operations per degree of freedom (DOF), and minimize memory accesses and global data movement. One way to describe the PA and MF algorithms is through the decomposition of a finite element operator into successive applications of simpler operators—, , and —that represent parallel communication, mesh topology, basis functions, and physics at quadrature points, respectively. This is illustrated in Figure 1.
@@ II-C Digital twin for tsunami early warning
Henneking et al. [undefu, undefv] recently developed a novel, scalable framework for real-time Bayesian inference for inverse problems governed by linear time-invariant dynamical systems. They applied this framework to create a physics-based digital twin for tsunami early warning in the Cascadia subduction zone—a 1000 km long megathrust fault located off the coast of the U.S. Pacific Northwest and Canada with the potential for a magnitude 8–9 earthquake and large-scale tsunami. Given real-time pressure data from seafloor sensors, the digital twin can infer the earthquake-induced seafloor motion and forecast tsunami wave heights in less than a second. This work won the 2025 ACM Gordon Bell Prize [undefl].
@@
The action of a finite element discretized PDE operator on a vector is typically a sparse operation with a FLOP count and memory access rate proportional to the number of DOF; for PDE-based applications, the finite element kernels performing this computation are therefore often the performance bottleneck of the entire application code. In other words, the highest potential for improving application performance is to increase the kernel throughput (which is the metric-of-interest that determines time-to-solution). In this paper, we demonstrate that kernel performance of high-order finite element discretizations can be significantly accelerated with the FP64 tensor cores. We focus on the performance improvement of a particular application example—the tsunami early warning digital twin. By using the application code and discretized operator presented in [undefl], we ensure that the application example, i.e. the particular PDE operator and discretization method, are application-driven rather than specifically chosen to enable the highest performance gains. Indeed, larger performance gains than the ones presented here could likely be achieved if the operators were discretized in such a way to best fit the tensor-core architecture, as will be discussed in Section III-E.
The tsunami dynamics are modeled by the acoustic–gravity wave equations that couple the propagation of ocean acoustic waves with surface gravity waves. This PDE model is derived by linearization of the conservation of mass and momentum around hydrostatic pressure in the compressible ocean, and the coupling to the surface gravity wave comes from a modified free surface boundary condition at the sea surface [undefw]. The model takes the form of a coupled first-order system in the velocity vector field , the pressure field , and the surface gravity wave height ,
@@ II-D CUDA and tensor cores
On NVIDIA GPUs, CUDA C++ employs a unique architecture called Single Instruction, Multiple Threads (SIMT) [undefx]. The same instruction is executed in groups of 32 threads (called a warp), and multiple warps of work are scheduled and pipelined on GPUs to achieve high compute throughput and memory bandwidth. The GPU hardware that executes the SIMT instructions is typically called CUDA cores. In a CUDA C++ kernel, multiple warps of threads are grouped into a thread block, also known as a cooperative thread array. Threads in the same thread block communicate via an on-chip, high-bandwidth, low-latency memory called shared memory [undefy].
@@
In addition to the CUDA cores, the latest NVIDIA GPUs also have tensor cores to speed up MMA instructions. Later in this work, we use the syntax // for matrix-matrix multiplications of shape -by- times -by- equals -by-. In this work, we use the FP64 MMA DMMA instruction that is first available in NVIDIA’s Ampere architecture.
@@ III Performance optimizations
In this section, we provide a detailed description of the performance analysis and optimization we performed for the finite element kernels from Section II. In Section III-A, we start with the original CUDA core kernels and their performance-limiting factors determined by Nsight Compute profiles. In Section III-B, we show how to use the FP64 tensor cores by invoking the DMMA instructions. Next, we describe the techniques we used to avoid shared memory bank conflicts in the FP64 tensor core computations in Section III-C and Section III-D. In Section III-E, we examine the performance-limiting factors of the DMMA kernels and finally discuss the fused PA and fused MF kernel optimizations in Section III-F.
@@ III-A The CUDA core kernels
A sequence of matrix-matrix multiplications are performed as part of the CUDA kernels that implement the actions of finite element discretized PDE operators on a vector (2). In the CUDA core kernels, for each matrix-matrix multiplication, each thread in the thread block calculates one element of output matrix . For each thread, the needed elements from input matrices and are loaded from shared memory, and the matrix elements are written back to shared memory. This results in the following skeleton CUDA code for such matrix-matrix multiplication:
@@
Note that the same input matrix elements in and are loaded by multiple threads, and there is no data sharing among the different threads after the shared memory loads.
@@
The total number of FP64 operations performed is
and the flop-over-byte ratio is
Therefore, shared memory bandwidth is the performance bottleneck of the kernel. This is confirmed by the outputs from the Nsight Compute profile shown in Table II: the metric “L1: Data Pipe Lsu Wavefronts” in the memory throughput breakdown hitting a percentage indicates that the kernel utilizes of the shared memory bandwidth.
@@ III-B Using tensor cores
of the DMMA instruction, shown as T[lane number]:c[0,1]. Quoted from [undefz].
@@
With the DMMA instruction, each warp in the thread block performs matrix-matrix multiplications collectively. Each thread of the warp loads a unique set of matrix elements from input matrix and : each matrix element in and is loaded only once among the threads in a warp. The DMMA instruction works in the following way [undefz]:
1.
2.
3.
Invoke the CUDA PTX instruction for DMMA with and from each thread as input [undefaa]. The CUDA PTX instruction can be seamlessly inserted into the CUDA C++ code with the inline PTX syntax.
4.
Fewer bytes are loaded from shared memory versus the CUDA core implementation because the input A/B matrix values are shared among the threads invoking the DMMA instruction in the same warp and only need to be loaded once per warp. To perform the target matrix-matrix multiplication with the DMMA instruction, the logical // indices of the DMMA instruction (//) are mapped to the target problem’s // indices (//):
@@
where is the warp index, and and are identical maps. can also be the following one:
@@ III-C Optimal mapping for bank conflict avoidance
For different shapes of matrix-matrix multiplications, different // can be chosen to avoid shared memory bank conflicts. For example, in order to perform an , , matrix-matrix multiplication, we use the four warps for the DMMA instruction with the following maps:
When each thread (also referred to as a lane in a warp) in a warp loads (stores) an 8-byte word from (to) the shared memory, the load (store) is done in two accessing phases. In the first phase, lanes 0–15 load (store) from (to) shared memory with each lane occupying two shared memory banks; in the second phase, lanes 16–31 do the same. As a result, if the first or the second half of the warp accesses 16 distinct shared memory banks during one 8-byte shared memory access, there is no shared memory bank conflict. With the maps shown in (5), each lane of a warp accesses one 8-byte FP64 number for a0, b0, c0 and c1. Figure 2 depicts the shared memory banks being accessed during the two access phases for matrices A, B and C. There are no shared memory bank conflicts for any of these accesses.
Apart from the (, , ) matrix shape shown above, we have also tuned the // for the following matrix shapes such that they are free from shared memory bank conflicts: , , , , and .
@@ III-E Tensor core kernel performance limiters
III-F PA and MF kernel optimizations
Our GPU kernel optimizations focused on the operations in (1), excluding the application of and , which are sparse matrix-vector products handling MPI communication. The finite element operator is mixed, i.e. the trial and test spaces are different, and so we need distinct and operators (and their respective transpose). By eliminating the initial from the preceding vector operation, the primary kernel comprises a sequence of two subkernels, and , such that , where
@@ IV-A Single-GPU performance
Figure 3 shows the throughput performance of the finite element kernels on NVIDIA’s GH200 Grace Hopper Superchip (left) and GB200 Grace Blackwell Superchip (right). For the various kernel types, the DMMA optimizations achieve 35% to 59% speedup over the original non-DMMA kernels in the saturated regime (i.e, for sufficiently many DOFs). Fusing loops into a single kernel and using tensor-core-enabled DMMA instructions (“DMMA Fused PA”) together yields a 2 performance gain compared to the original PA kernel.
@@
Table VI details the throughput performance (GDOF/s) and the performance per Watt (MDOF/W) on both GH200 and GB200 for the original and the DMMA kernels, for a problem of 540 million DOF. In addition to higher performance, the DMMA kernels also achieve higher performance per Watt than the original kernels. Usage of FP64 tensor cores improves performance per Watt by 18% on GB200 and 27% on GH200. When fusion is added, performance per Watt improves overall by 72% on GB200 and 83% on GH200. We note that GB200 achieves a lower performance per Watt than GH200 for each of the kernel types. The reasons for this include that (1) GB200 has higher idle GPU power, (2) GB200 runs at a 4% higher clock frequency which, while adding additional performance, can lead to less efficient computations, and (3) GB200 has much higher memory bandwidth and tensor core performance for lower precisions (FP16, BF16, FP8, FP6 and FP4), while the kernels of interest here do not make high utilizations of any of these compute resources.
@@ IV-B Multi-GPU performance
Scalability of the MFEM implementation was tested on CSCS’s Alps system. Alps is an HPE Cray EX supercomputer with 2,688 nodes, each containing 4 NVIDIA GH200 Superchips. Each GH200 integrates a Grace CPU (72 Arm cores) and an H100 GPU (96 GB HBM3, 34 TFLOP/s FP64 and 67 TFLOP/s FP64 tensor core peak throughput). The system is connected by an HPE Slingshot-11 dragonfly interconnect with at most three network hops between any two nodes. The theoretical system peak is 574.8 PFLOP/s; and the system achieved a peak of 434.9 PFLOP/s securing the #8 spot in the November 2025 TOP500 list [undefab].
@@
The strong scalability results, shown in Figure 4, demonstrate ideal (linear) strong scaling for all kernels over the first 8 increase of nodes (from 144 to 1,152 GH200); at full system size (i.e., over a 64 increase of nodes), the finite element implementations achieve excellent 86–91% strong parallel efficiencies. Weak scalability results, depicted in Figure 5, show that ideal (linear) weak scaling (i.e, 100% weak parallel efficiency) is achieved for all of the six kernel implementations over a 64 increase of nodes.

### 09 Apr 2026 v2

@@ I Introduction
Starting in 2020 with the Ampere A100 GPU, NVIDIA introduced FP64 tensor cores supporting native IEEE double precision MMA (DMMA) instructions. For double precision calculations, these tensor cores deliver a 2 increase in peak floating-point rates [undefd]. In recent years, other vendors have introduced tensor cores, or analogous concepts, in their compute architectures. In this paper, we focus specifically on the programming of tensor cores on two of NVIDIA’s state-of-the-art chips: the Grace Hopper GH200 Superchip [undefe] and the Grace Blackwell GB200 Superchip [undefb].
@@
We build on this prior work and apply tensor cores to a full finite element production application—a digital twin for tsunami early warning—that is already well-optimized and won the Gordon Bell Prize in 2025 [undefl]. The application code solves a coupled acoustic–gravity wave propagation problem as a first-order system of partial differential equations (PDEs) discretized with high-order finite elements. Specifically, we decompose the tensor contractions into small GEMMs of order and program the FP64 tensor cores directly in CUDA kernels. We use Parallel Thread Execution (PTX) instructions to directly program tensor cores on state-of-the-art chips like NVIDIA’s Grace Hopper GH200 and Grace Blackwell GB200. We also create custom mappings to handle irregular matrix shapes and fuse kernels for additional performance. This leads to significant performance (up to 2) and energy efficiency (up to 83%) gains in a complex, already well-tuned application code. The resulting application demonstrates near-ideal weak scaling and 90% strong scaling efficiency on 9,216 GPUs.
@@
To the best of our knowledge, the first example of using directly programmed FP64 tensor cores in a complex, PDE-based HPC application, in order to improve throughput of high-order finite element kernels (by up to 59%). We show that shared memory data motion, not FLOPs, is the performance bottleneck, and demonstrate strong correlation between tensor core data motion reduction and improved application performance.
@@
Significant finite element kernel performance gains (up to 2) and energy efficiency gains (up to 83%) by combining loop fusion techniques with tensor core kernels.
Demonstration of exascale full-system scalability of the application code for each of the implemented kernel versions on a leading-edge supercomputer.
The rest of the paper is organized as follows. Section II provides a brief overview of the MFEM library, the finite element algorithms and application used, and the NVIDIA FP64 DMMA tensor cores. We then discuss how we optimized the key finite element kernels in Section III. Section IV shows the performance and energy efficiency improvements from our optimizations. In Section V and Section VI, respectively, we describe related work and summarize our conclusions.
@@ II-A MFEM library
II-B Finite element assembly algorithms on GPUs
@@
GPU-accelerated PA kernels cover a broad range of operators, enabling end-to-end acceleration of matrix-free solvers through low-order-refined preconditioning [undeft]. Kernel fusion methods enhance strong scalability and allow attaining peak performance on problem sizes 5–10 times smaller than conventional methods. These matrix-free approaches, distinct from traditional sparse matrix assembly, significantly reduce memory footprint by avoiding global matrix storage, enhance arithmetic intensity to achieve near-optimal floating-point operations per degree of freedom (DOF), and minimize memory accesses and global data movement.
One way to describe the PA and MF algorithms is through the decomposition of a finite element operator into successive applications of simpler operators—, , and —that represent parallel communication, mesh topology, basis functions, and physics at quadrature points, respectively. This is illustrated in Figure 1.
@@ II-C Digital twin for tsunami early warning
Henneking et al. [undefu, undefv] recently developed a novel, scalable framework for real-time Bayesian inference for inverse problems governed by linear time-invariant dynamical systems. They applied this framework to create a physics-based digital twin for tsunami early warning in the Cascadia subduction zone [undefw]—a 1000 km long megathrust fault located off the coast of the U.S. Pacific Northwest and Canada with the potential for a magnitude 8–9 earthquake and large-scale tsunami. Given real-time pressure data from seafloor sensors, the digital twin can infer the earthquake-induced seafloor motion and forecast tsunami wave heights in less than a second. This work won the 2025 ACM Gordon Bell Prize [undefl].
@@
The action of a finite element discretized PDE operator on a vector is typically a sparse operation with a FLOP count and memory access rate proportional to the number of DOF; for PDE-based applications, the finite element kernels performing this computation are therefore often the performance bottleneck of the entire application code. In other words, the highest potential for improving application performance is to increase the kernel throughput (which is the metric-of-interest that determines time-to-solution). In this paper, we demonstrate that kernel performance of high-order finite element discretizations can be significantly accelerated with the FP64 tensor cores. We focus on the performance improvement of a particular application example—the tsunami early warning digital twin. By using the application code and discretized operator presented in [undefl], we ensure that the application example, i.e., the particular PDE operator and discretization method, is application-driven rather than specifically chosen to enable the highest performance gains. Indeed, larger performance gains than the ones presented here could likely be achieved if the operators were discretized in such a way to best fit the tensor-core architecture, as will be discussed in Section III-E.
The tsunami dynamics are modeled by the acoustic–gravity wave equations that couple the propagation of ocean acoustic waves with surface gravity waves. This PDE model is derived by linearization of the conservation of mass and momentum around hydrostatic pressure in the compressible ocean, and the coupling to the surface gravity wave comes from a modified free surface boundary condition at the sea surface [undefx]. The model takes the form of a coupled first-order system in the velocity vector field , the pressure field , and the surface gravity wave height ,
@@ II-D CUDA and tensor cores
On NVIDIA GPUs, CUDA C++ employs a unique architecture called Single Instruction, Multiple Threads (SIMT) [undefy]. The same instruction is executed in groups of 32 threads (called a warp), and multiple warps of work are scheduled and pipelined on GPUs to achieve high compute throughput and memory bandwidth. The GPU hardware that executes the SIMT instructions is typically called CUDA cores. In a CUDA C++ kernel, multiple warps of threads are grouped into a thread block, also known as a cooperative thread array. Threads in the same thread block communicate via an on-chip, high-bandwidth, low-latency memory called shared memory [undefz].
@@
In addition to the CUDA cores, the latest NVIDIA GPUs also have tensor cores to speed up MMA instructions. Later in this work, we use the syntax // for matrix-matrix multiplications of shape -by- times -by- equals -by-. In this work, we use the FP64 DMMA instruction that is first available in NVIDIA’s Ampere architecture.
@@ III Performance optimizations
In this section, we provide a detailed description of the performance analysis and optimization we performed for the finite element kernels from Section II. In Section III-A, we start with the original CUDA core kernels and their performance-limiting factors determined by Nsight Compute profiles. In Section III-B, we show how to use the FP64 tensor cores by invoking the DMMA instructions. Next, we describe the techniques we used to avoid shared memory bank conflicts in the FP64 tensor core computations in Section III-C and Section III-D. In Section III-E, we examine the performance-limiting factors of the DMMA kernels. In Section III-F, we explain why the CUDA PTX DMMA instruction is used over other frameworks, and finally discuss the fused PA and fused MF kernel optimizations in Section III-G.
@@ III-A The CUDA core kernels
A sequence of matrix-matrix multiplications is performed as part of the CUDA kernels that implement the actions of finite element discretized PDE operators on a vector (2). The kernels used in the Gordon Bell Prize-winning work [undefl] use the CUDA cores to perform these matrix-matrix multiplications. In these CUDA core kernels, for each matrix-matrix multiplication, each thread in the thread block calculates one element of output matrix . For each thread, the needed elements from input matrices and are loaded from shared memory, and the matrix elements are written back to shared memory. This results in the following skeleton CUDA code for such matrix-matrix multiplication:
@@
Note that the same input matrix elements in and are loaded by multiple threads, and there is no data sharing among the different threads after the shared memory loads. The total number of FP64 operations performed is
and the FLOP-over-byte ratio is
Therefore, it is the shared memory bandwidth, not the FP64 compute FLOP/s, that limits the performance of the kernel. This is confirmed by the outputs from the Nsight Compute profile shown in Table II: the metric “L1: Data Pipe Lsu Wavefronts” in the memory throughput breakdown hitting a percentage indicates that the kernel utilizes of the shared memory bandwidth. In contrast, in the same table, the metric “SM: Pipe Fp64 Cycles Active” hits only .
@@ III-B Using tensor cores
With the DMMA instruction, each warp in the thread block performs matrix-matrix multiplications collectively. Each thread of the warp loads a unique set of matrix elements from input matrices and : each matrix element in and is loaded only once among the threads in a warp. The DMMA instruction works in the following way [undefaa]:
1.
2.
3.
Invoke the CUDA PTX instruction for DMMA with and from each thread as input [undefab]. The CUDA PTX instruction can be seamlessly inserted into the CUDA C++ code with the inline PTX syntax.
4.
This results in the following skeleton CUDA code for performing matrix-matrix multiplications with DMMA tensor cores with the CUDA PTX DMMA instruction:
Fewer bytes are loaded from shared memory versus the CUDA core implementation, because the input A/B matrix values are shared among the threads invoking the DMMA instruction in the same warp and only need to be loaded once per warp. To perform the target matrix-matrix multiplication with the DMMA instruction, the logical // indices of the DMMA instruction (//) are mapped to the target problem’s // indices (//):
@@
of the DMMA instruction, shown as T[lane number]:c[0,1]. Quoted from [undefaa].
@@
where is the warp index, and and are identity maps. can also be the following one:
@@ III-C Optimal mapping for bank conflict avoidance
For different shapes of matrix-matrix multiplications, different // can be chosen to avoid shared memory bank conflicts. For example, in order to perform an , , matrix-matrix multiplication, we use four warps for the DMMA instruction with the following maps:
When each thread (also referred to as a lane in a warp) in a warp loads (stores) an 8-byte word from (to) the shared memory, the load (store) is done in two accessing phases. In the first phase, lanes 0–15 load (store) from (to) shared memory with each lane occupying two shared memory banks; in the second phase, lanes 16–31 do the same. As a result, if the first or the second half of the warp accesses 16 distinct shared memory banks during one 8-byte shared memory access, there is no shared memory bank conflict. With the maps shown in (5), each lane of a warp accesses one 8-byte FP64 number for a0, b0, c0 and c1.
Figure 2 depicts the shared memory banks being accessed during the two access phases for matrices A, B and C. There are no shared memory bank conflicts for any of these accesses. Apart from the () matrix shape shown above, we have also tuned the // for the following matrix shapes such that they are free from shared memory bank conflicts: , , , , and .
@@ III-E Tensor core kernel performance limiters
III-F Choosing CUDA DMMA instruction over other frameworks
In this work, the CUDA PTX DMMA instruction is chosen over other CUDA libraries like CUTLASS [undefac] or CUBLAS [undefad], because our small matrices—such as —require precise control for custom thread-to-fragment mappings (5) and bank conflict elimination (Figure 2). CUTLASS is optimized for larger GEMMs and lacks the needed granularity; its primary method to avoid shared memory bank conflicts is swizzling, which requires much larger alignments (e.g., 128 bytes) and shapes (e.g., ). While padding could enable the use of CUTLASS kernel-level interfaces, with the small matrix sizes we are targeting, the majority of the shared memory bandwidth and FLOP/s would be wasted on the padding, rendering CUTLASS much less efficient than our approach in this case.
CUBLAS, on the other hand, does not support such unconventional scenarios with small, irregular GEMMs and kernel fusions. CUBLASDx [undefae], which is still under development, addresses our needs and may in the future provide an efficient alternative to our approach.
Finally, we note that invoking the CUDA PTX DMMA instruction with the inline PTX syntax, as previously shown in the skeleton CUDA code for DMMA tensor cores, requires just about the same amount of developer effort as invoking other CUDA device intrinsic functions when writing a CUDA C++ kernel.
III-G PA and MF kernel optimizations
Our GPU kernel optimizations focused on the operations in (1), excluding the application of and , which are sparse matrix-vector products handling MPI communication. The finite element operator is mixed, i.e., the trial and test spaces are different, and so we need distinct and operators (and their respective transposes).
By eliminating the initial from the preceding vector operation, the primary kernel comprises a sequence of two subkernels, and , such that , where
@@ IV-A Single-GPU performance
Figure 3 shows the throughput performance of the finite element kernels on NVIDIA’s GH200 Grace Hopper Superchip (left) and GB200 Grace Blackwell Superchip (right). For the various kernel types, the DMMA optimizations achieve 35% to 59% speedup over the original non-DMMA kernels in the saturated regime (i.e., for sufficiently many DOFs). Fusing loops into a single kernel and using tensor-core-enabled DMMA instructions (“DMMA Fused PA”) together yields a 2 performance gain compared to the original PA kernel.
@@
Table VI details the throughput performance (GDOF/s) and the performance per Watt (MDOF/W) on both GH200 and GB200 for the original and the DMMA kernels, for a problem of 540 million DOF. In addition to higher performance, the DMMA kernels also achieve higher performance per Watt than the original kernels. Usage of FP64 tensor cores improves performance per Watt by 18% on GB200 and 27% on GH200. When fusion is added, performance per Watt improves overall by 72% on GB200 and 83% on GH200. We note that GB200 achieves a lower performance per Watt than GH200 for each of the kernel types. The reasons for this include: (1) GB200 has higher idle GPU power, (2) GB200 runs at a 4% higher clock frequency which, while adding additional performance, can lead to less efficient computations, and (3) GB200 has much higher memory bandwidth and tensor core performance for lower precisions (FP16, BF16, FP8, FP6 and FP4), while the kernels of interest here do not make high utilizations of any of these compute resources.
@@ IV-B Multi-GPU performance
Scalability of the MFEM implementation was tested on CSCS’s Alps system. Alps is an HPE Cray EX supercomputer with 2,688 nodes, each containing 4 NVIDIA GH200 Superchips. Each GH200 integrates a Grace CPU (72 Arm cores) and an H100 GPU (96 GB HBM3, 34 TFLOP/s FP64 and 67 TFLOP/s FP64 tensor core peak throughput). The system is connected by an HPE Slingshot-11 dragonfly interconnect with at most three network hops between any two nodes. The theoretical system peak is 574.8 PFLOP/s; and the system achieved a peak of 434.9 PFLOP/s securing the #8 spot in the November 2025 TOP500 list [undefaf].
@@
The strong scalability results, shown in Figure 4, demonstrate ideal (linear) strong scaling for all kernels over the first 8 increase of nodes (from 144 to 1,152 GH200); at full system size (i.e., over a 64 increase of nodes), the finite element implementations achieve excellent 86–91% strong parallel efficiencies. Weak scalability results, depicted in Figure 5, show that ideal (linear) weak scaling (i.e., 100% weak parallel efficiency) is achieved for all of the six kernel implementations over a 64 increase of nodes.